Grade upto college level Magnetism

Q. A circular coil A of 100 turns and radius 2 cm, is placed co-centrically inside another circular coil of B of 1000 turns of radius 20 cm, both being in the same plane. When the current in the coil B decreases from 5A to 3A in 0.04s, the induced enf in the coil A is : a) 1.97 mV b) 2.53mV c) 1.53mV d)1.06mV


Plz expail it elaborately.

Shane Macguire
12 Years agoGrade upto college level

1 Answer

ROSHAN MUJEEB
6 Years ago

Solution :
N1.μN2i2r2.πr21=MiN1.μN2i2r2.πr12=Mi
M=μN2i.πr212r2M=μN2i.πr122r2
=4π×10−7×100×100×π(0.02)22×0.2=4π×10-7×100×100×π(0.02)22×0.2
=4×10−5H=4×10-5H
(b) Induced emf inpp:
ep=Mdi2dr=4×10−5×20.04ep=Mdi2dr=4×10-5×20.04
=2×10−3V=2×10-3V